12.5 ml of 20x EDTA were diluted with 237.5ml of H2O. What is the final conc?
Answer #1

we use the relation: Volume * Concentration = constant, here we use initial conc. 20x, Volume = 12.5. Final concentration M and Final volume = 12.5 + 237.5 = 250 mL

12.5 * 20x = 250 * M

SOlving, M = 1x solution of EDTA